OpenStudy (anonymous):

Which description is true about the expression 12+2a+3(b)-6? 4 terms 1 variable term 1 coefficient 3 constant terms

OpenStudy (anonymous):

none of them

OpenStudy (anonymous):

4 terms.

OpenStudy (anonymous):

D is probable abswer

OpenStudy (anonymous):

4 terms

OpenStudy (anonymous):

no

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

there are 2 constants -_-

OpenStudy (anonymous):

12+2a+3-6=2a+3b+6

OpenStudy (anonymous):

No how can it have 3 constants when only 12 -6 = 6 (1 constant)

OpenStudy (anonymous):

there are three constant terms 2,3,6

OpenStudy (anonymous):

2 and 3 are coefficient of a and b which is constant

OpenStudy (anonymous):

3b is not a constant

OpenStudy (anonymous):

@itsmylife... No, it's not... Any equation has only one constant (the term without a variable)

OpenStudy (anonymous):

and you don't need to modify the expression

OpenStudy (anonymous):

@sparik1997 3 is constant i am not talking about 3b

OpenStudy (anonymous):

consider it as it is!

OpenStudy (anonymous):

A constant is a term WITHOUT a variable!

OpenStudy (anonymous):

@ayushprasad look if D is not the answer then none of the choices are correct and you need to see constants not variables attached to them

OpenStudy (anonymous):

it has got only 2 variables a and b

OpenStudy (anonymous):

I think 4 terms is right, I will find out once I get it graded. Lol.

OpenStudy (anonymous):

OMG... those are COEFFECIANTS!!! NOT CONSTANTS!!!

OpenStudy (anonymous):

and 12-6=6 @ayushprasad

OpenStudy (anonymous):

You don't say?

OpenStudy (anonymous):

i know but i am saying in the whole equation there are 3 terms 2a+3b+6 =0 and two variables

OpenStudy (anonymous):

no... there are 2 variables and 2 constants. let's see

OpenStudy (anonymous):

okay look go with what heart says, make a guess and proceed

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

Okay, since you are being stubborn, I'll just quit and move on to another question. Bye!

OpenStudy (anonymous):

alright i know you're taking it as coefficient and i am taking it as constant because i see none of the options as correct one

OpenStudy (anonymous):

see ya

OpenStudy (anonymous):

Btw, who's the one with the medal? This guy :D

OpenStudy (anonymous):

lol i donno

OpenStudy (anonymous):

You mad bro?

OpenStudy (anonymous):

i am ;)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@because100 what was the answer?

OpenStudy (anonymous):

I still have 12 more questions to do before I get the right answer.

OpenStudy (anonymous):

ok, tell us when you are finished

OpenStudy (anonymous):

Will do.

OpenStudy (anonymous):

4 terms was correct.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Lol ^_^ Thanks for answering.

OpenStudy (anonymous):

;)